OpenStudy (anonymous):

Help please :) ? If f(x) = 2x2 (x + 1) then f ′(1) is equal to i get everything but 10 and 10 is the answer :)

OpenStudy (anonymous):

\[ f(x)= 2x^2(x+1) \]?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

If f(x) = 2x^2 times (x + 1) ...? so f(x) = 2x^2 (x + 1) = 2x^3 + 2x^2 so f'(x) = 6x^2 + 4x ... yeah?

OpenStudy (anonymous):

What is your work?

OpenStudy (jack1):

then sub in x = 1

OpenStudy (jack1):

... =10

OpenStudy (anonymous):

i get what you did but is it okay if you explain the part after the so it equals like there is a 3 to the power of instead of a a 2 why did you change it? @Jack1

OpenStudy (anonymous):

ohhhhhh ok thanks :)

OpenStudy (jack1):

cool, u got that all clear now?

OpenStudy (jack1):

sorry, on another q, too slow to come back

OpenStudy (anonymous):

yes thanks :)

OpenStudy (jack1):

sáll good, happy trails ;D